# Identity-Obscured Neural Radiance Fields: Privacy-Preserving 3D Facial Reconstruction

### v1

@@ Abstract
Neural radiance fields (NeRF) typically require a complete set of images taken from multiple camera perspectives to accurately reconstruct geometric details. However, this approach raise significant privacy concerns in the context of facial reconstruction. The critical need for privacy protection often leads invidividuals to be reluctant in sharing their facial images, due to fears of potential misuse or security risks. Addressing these concerns, we propose a method that leverages privacy-preserving images for reconstructing 3D head geometry within the NeRF framework. Our method stands apart from traditional facial reconstruction techniques as it does not depend on RGB information from images containing sensitive facial data. Instead, it effectively generates plausible facial geometry using a series of identity-obscured inputs, thereby protecting facial privacy.
1 Introduction
3D facial models play a crucial role in various human-related applications, particularly in identity authentication and recognition systems within financial and security contexts. Nowadays, the emerging techniques like Neural Radiance Fields (NeRF) [18] have significantly enhanced the accuracy of 3D facial reconstruction. Nevertheless, traditional methods for neural surface reconstruction depend heavily on acquiring 3D facial data. This often involves collecting images of individuals, specifically those capturing detailed geometric and color information from the front view. This data collection process may give rise to privacy concerns if the collected facial data is unintentionally disclosed.
The development of generative models exacerbates the adverse consequences of front-face image leakage, as they can generate a large number of realistic privacy-compromising face images using only a minimal set of leaked face photos. For instance, Vendrow et al. propose a GAN [30] to create facial images convincingly real to a victim by utilizing leaked front-face photos from a database. Shahreza et al. [29] go a step further by using GNeRF to generate multi-view realistic fake face images of a victim from real face photos of an identity. These generated images have the capability to deceive some state-of-the-art face recognition (FR) systems through adversarial attacks. Yu et al. [43] introduce NOFA, reconstructing high-fidelity facial avatars with one-shot learning on a single facial image. The potential risks become significantly more perilous when data, particularly sensitive front-face images, falls into malicious hands. Therefore, it is crucial to reevaluate the reconstruction problem from a privacy perspective, acknowledging that facial privacy should ideally be safeguarded in the context of reconstruction tasks. Hence, there is an urgent need to explore specialized 3D facial reconstruction techniques that prioritize front-face privacy protection.
Figure 1: Comparison of Paradim with other facial reconstruction methodologies: (a) VolSDF [41] can recover geometric details but lacks privacy considerations. (b) mm3DFace [36] and similar methods can protect privacy but sacrifice geometric details. (c) Our method excels in recovering fine geometric details while preserving privacy.
This paper presents a comprehensive end-to-end strategy that establishes a privacy protection framework and efficiently utilizes both privacy-neutral and privacy-preserving data in the reconstruction process. Privacy exposure primarily results from the interception of RGB information in facial images. Nevertheless, we also observe that areas with rapidly changing RGB values in the image likely correspond to regions with significant geometric variations. Consequently, by eliminating the RGB information and preserving solely the color variation amplitude, we can reconstruct plausible geometry details while maintaining privacy.
As illustrated in Figure 1, we are the first to accomplish geometric detail reconstruction while ensuring privacy. Specifically, we conceal frontal face information, making the processed face visually indistinguishable and irretrievable. Importantly, accurate geometry reconstruction remains possible. In Figure 1 (a), Neural Surface Reconstruction demonstrates the restoration of excellent geometric details but does not address the specific privacy requirements in face reconstruction processes. Other face reconstruction approaches, such as those utilizing radar in Figure 1 (b) or infrared camera technology, may meet privacy requirements but fall short in reconstructing intricate geometric details. Our method classifies input data and employs staged processing and reconstruction, enabling the restoration of geometric details while adhering to privacy considerations.
@@ 1 Introduction
In this work, we achieve a balance between face reconstruction demands and user privacy concerns, offering a solution for reliable and secure facial geometry generation across various applications. Our main contributions are summarized as:
  • Our paper stands out as the first to tackle the crucial issue of privacy protection in NeRF-based reconstruction. We not only introduce a novel perspective on privacy concerns but also provide a practical and pioneering solution.
  • We propose a generic and irreversible processing method that can make images with sensitive information blurry and devoid of noticeable color details.
  • By incorporating information from privacy-preserving and privacy-neutral images into the neural radiance fields pipeline, we achieve a high level of geometric detail.
2 Related Work
@@
3D Morphable Models (3DMM) [2] leverage principal component analysis to represent facial geometry and appearance. However, this method falls short in capturing details such as wrinkles, the interior of the mouth, and hair, so it may not fully satisfy appearance requirements. i3DMM [42] introduces an implicit function to model both the geometry and appearance of the human head with various attributes like shape, expression, and hairstyle. It’s important to note that, similar to 3DMM, both approaches rely on 3D data for their modeling. Recent advancements in Neural Radiance Fields (NeRF) [18, 1] have excelled in novel view synthesis due to their compact and powerful representation capability, relying solely on a set of multi-view images. Consequently, numerous works [9, 24] have yielded detailed geometry [46, 37] and achieved a photo-realistic appearance [45, 12] in modeling human heads.
@@
Sign distance fields (SDF) [23] and occupancy fields [17], showcase their representative ability over explicit representations, i.e., mesh and point cloud. DVR [21] and IDR [40] focus on differentiating the surface rendering pipeline based on multi-view images. They incorporate corresponding masks to distinguish objects from the background during the training process. NeuS [31] and VolSDF [41] refine the geometry reconstruction of NeRF by introducing a scheme that converts SDF to density, enhancing the surface representation in NeRF. Recent approaches [26, 33, 34] combine volume rendering with multi-scale hashing in Instant-NGP [20] and displacement fields to learn detailed surfaces through implicit functions from sets of multiview images.
@@
We summarize three primary approaches. The first involves an algorithm based on adversarial generation networks, aimed at deceiving unauthorized facial recognition by generating fake images highly similar to the original ones [15, 19]. However, this method is primarily targeted at public social platforms [5] or specific facial recognition (FR) systems that provide data for learning [4]. Given that the server is monitored and not randomly accessed, this form of privacy protection may not be convincing enough for users who might opt to upload visually unfriendly images. The second category involves the use of cryptographic techniques. Cryptographic techniques, including homomorphic encryption [7, 10, 27], secure multiparty computation [16, 35, 39], and other encryption primitives [13], are used to encrypt original images securely. This approach ensures the secure identification of encrypted images while maintaining high confidentiality. However, it may introduce higher latency and computational costs. The third group of methods focuses on obfuscation techniques. These methods encompass actions such as implementing blurring [11], introducing noise [44], applying masking [28], utilizing filtering [47], and employing image transformation [32]. Obfuscation methods are regarded as more practical and are extensively utilized for safeguarding facial privacy. Despite their widespread use due to operational feasibility, it’s important to note that the decrease in the quality of facial images is irreversible. Without a well-designed complete end-to-end process, there is a risk of failure in subsequent downstream tasks.
3 Method
Figure 2: Pipeline. We simulate the data flow process where our model reconstructs the human head geometry, particularly the face, from identity-obscured input. Users have the option to select photos that require privacy protection through a general operator denoted by $\sigma$. Following this, both privacy-neutral and privacy-preserving images undergo processing and are sent to the server for geometric reconstruction. The server-side reconstruction occurs in two stages. In the first stage, we employ privacy-neutral images to establish the foundational geometry. In the second stage, privacy-preserving images are used for further refinement of facial geometry.
3.1 Preliminaries
Neural volume rendering. As demonstrated by NeRF [18], it characterizes a 3D scene by employing volume density and color fields. Given known camera poses and ray directions, we conduct sampling along the rays, predicting both the color $\mathbf{c}_{i}$ and density $\sigma_{i}$ at each sample point $\mathbf{x}_{i}$ using Multi-Layer Perceptron (MLPs). The volume rendering process entails the integrating color radiance across the sampled points along the ray. Consequently, we approximate the rendered color for a specific pixel as:
@@ 3.1 Preliminaries
Recent contributions have been centered around neural volume rendering based on SDF representation [31, 41]. These methods utilize MLPs to implicitly represent a 3D scene by predicting the SDF and color $\mathbf{c}_{i}$ relative to the viewpoint. This differs from the initial approach presented in [18], which focused on predicting color and density. The extraction of the zero-level surface of the SDF in these methods results in a more reasonable geometric representation, showcasing significant effectiveness in reconstructing objects with smooth geometry.
@@
3.2 Two-stage Training
In our assumptions, server-stored data should be anonymized and non-identifiable. When processing facial data and generating geometric representations, the server should neither access nor alter sensitive user facial images. This crucial requirement must be met by our algorithm without compromising the practical utility of geometric reconstruction. To achieve this, we have implemented a categorization system for the images submitted by users, classifying them into two main categories: privacy-neutral and privacy-preserving. Privacy-neutral pictures are employed directly by the server during the neural surface reconstruction training process, whereas images that necessitate privacy protection undergo processing through our preserving operator. These images cannot be directly utilized for reconstruction in the second stage. Instead, they will undergo specialized processing to align with our reconstruction objectives.
In our proposed method, we employ a two-stage training framework for geometric reconstruction in Figure 2. In the first stage, we train a neural radiance field using privacy-neutral images, such as images taken of the back of the head. This initial privacy-neutral training stage enables the reconstruction of essential low-frequency information. The success of this stage depends on the quantity of privacy-neutral data users provide. While this information might have been deemed less valuable for facial geometry learning in the past, it proves to be beneficial in our method. Unless unavoidable circumstances require the use of templates, as discussed in Section 3.4, the privacy-neutral information provided by the user can be fully utilized. This enhances the geometric accuracy in the first stage of reconstruction, bringing it closer to the user’s geometric features, ultimately contributing to the overall quality of the reconstruction in the second stage.
In the second stage, we utilize privacy-preserving gradient image data uploaded by users for the reconstruction process. For detailed information on image privacy protection, refer to Section 3.3. In this stage, our supervision focuses solely on color variation information present in the original images, unlike the RGB color information used in neural radiance fields.
While the privacy-preserving images appear visually unfriendly and blurry, they contain valuable color variation details crucial for refining geometric intricacies. Building on the foundation established in the first stage, we train for new frontal face viewpoints. We compute gradients of the rendered images, transforming them into multi-view color gradient modulus through a general operator $\sigma$. These modulus are then compared to the facial privacy-preserving gradient information obtained from the user after passing through the operator. To optimize this stage, we introduce a novel loss equation, the gradient loss $\mathcal{L}_{\text{grad}}$, to guide the learning of geometric information in the privacy-preserving stage:

### v2

# Do Not DeepFake Me: Privacy-Preserving Neural 3D Head Reconstruction Without Sensitive Images

@@ Abstract
While 3D head reconstruction is widely used for modeling, existing neural reconstruction approaches rely on high-resolution multi-view images, posing notable privacy issues. Individuals are particularly sensitive to facial features, and facial image leakage can lead to many malicious activities, such as unauthorized tracking and deepfake. In contrast, geometric data is less susceptible to misuse due to its complex processing requirements, and absence of facial texture features. In this paper, we propose a novel two-stage 3D facial reconstruction method aimed at avoiding exposure to sensitive facial information while preserving detailed geometric accuracy. Our approach first uses non-sensitive rear-head images for initial geometry and then refines this geometry using processed privacy-removed gradient images. Extensive experiments show that the resulting geometry is comparable to methods using full images, while the process is resistant to DeepFake applications and facial recognition (FR) systems, thereby proving its effectiveness in privacy protection.
Introduction
In the rapidly evolving digital era, 3D facial reconstruction technology has become an indispensable component in fields such as modeling, virtual reality, and digital entertainment. Recent advancements, particularly the emergence of techniques like Neural Radiance Fields (NeRF) (Mildenhall et al. 2020) and Signed Distance Function (SDF)-based methods (NeuS/VolSDF) (Wang et al. 2021; Yariv et al. 2021), have significantly enhanced the accuracy of reconstruction, heralding a revolutionary progression in this domain. However, these methods often depend on multi-view, high-resolution facial images. While such images provide substantial convenience and precision in reconstruction, they raise significant privacy concerns due to the sensitive nature of the detailed facial information they contain. This reliance on sensitive images inevitably compromises privacy.
Extensive research has consistently demonstrated that individuals exhibit a markedly higher sensitivity to facial features compared to other body parts (Zebrowitz and Montepare 2008). Even minor facial imperfections, such as acne, scars, or uneven skin tone, can exert considerable psychological effects (Hamler et al. 2022; Mekeres et al. 2023). The potential misuse of facial information raises substantial concerns, including risks of social discrimination, psychological distress, and other negative consequences (Abrams et al. 2020). Furthermore, the necessity for protecting facial privacy extends beyond individual sensitivities to include the broader risks associated with the leakage of facial images (Ciftci, Yuksek, and Demir 2023). Current image-based facial recognition (FR) systems have reached a level of sophistication that enables the extraction and analysis of extensive information from facial images (Kortli et al. 2020). In the event of facial image leakage, this data could be exploited for unauthorized digital tracking, surveillance, or other malicious purposes (Hill 2022). Moreover, facial images provide the precise details necessary for creating DeepFake (Chadha et al. 2021), further amplifying privacy risks.
In contrast, geometric data is generally less sensitive compared to facial images. Geometric data alone lacks the detailed facial textures necessary for impersonation and forgery, such as those used in DeepFake technology (Westerlund 2019). This absence of detailed textures significantly diminishes the risk of misuse. Additionally, collecting, processing, and analyzing 3D data is more complex and resource-intensive (Uy et al. 2021), which hinders efficient tracking and recognition. The technology for handling 3D data is less developed and less widespread than that for 2D images (Guo et al. 2023), further reducing the risk of unauthorized monitoring and tracking. Thus, while facial geometry also reveals some information, it is much harder to exploit in DeepFake and FR systems, resulting in fewer malicious effects than the facial images used in current face reconstruction methods. This phenomenon highlights the urgent need to improve current methods and explore high-quality 3D reconstruction techniques that do not depend on high-risk facial images.
Existing privacy protection methods, such as image blurring (Jiang et al. 2023) or facial masking (Sun et al. 2018), while somewhat effective in 2D image processing, face numerous challenges in 3D reconstruction tasks. mm3DFace (Xie et al. 2023) is one of the approaches that address facial privacy in 3D reconstruction by using mmWave signals to extract geometric features, thus avoiding facial images and offering some level of privacy protection. It can track 68 facial landmarks, which is useful for expression analysis. However, it cannot capture geometric details which is necessary for nuanced applications. Currently, there is a lack of a comprehensive end-to-end strategy capable of reconstructing detailed facial geometry without relying on sensitive facial images.
To address this challenge, we propose an innovative 3D neural reconstruction method without dependence on sensitive facial images, representing the first systematic attempt to address privacy protection issues in the field of neural surface reconstruction. Our method employs a two-stage reconstruction process using two types of images: privacy-neutral images (e.q. rear-head images) and privacy-protected images, which are initially facial images processed to remove sensitive information. We first utilize privacy-neutral images to establish the basic geometric structure and then use privacy-protected images to refine and perfect the 3D reconstruction. This approach completely eliminates the reliance on sensitive facial RGB images while still enabling the reconstruction of stable and detailed geometric head models, thus avoiding the risk of sensitive image leakage.
@@ Introduction
Our research strikes a balance between the requirements of head reconstruction and user privacy protection, offering a reliable and secure head geometry reconstruction solution for various application scenarios. Our main contributions can be summarized as follows:
  • We introduce a novel end-to-end neural reconstruction pipeline that effectively protects facial privacy in 3D modeling. Our method eliminates the exposure of sensitive information while maintaining high-quality geometric reconstruction, thereby broadening NeRF’s applications in privacy-sensitive scenarios and laying a foundation for privacy-protected advancements.
  • We develop a unique two-stage training process that balances reconstruction stability and detail fidelity. This approach first demonstrates that gradient images can effectively contribute to geometric reconstruction, preventing reliance on sensitive facial images.
  • Our method significantly improves the security and effectiveness of neural reconstruction in sensitive contexts. Using images that lack detailed facial textures, protects facial privacy and renders them ineffective for common exploitation techniques such as DeepFake creation and facial recognition.
Related Work
@@
3D Morphable Models (3DMM) (Blanz and Vetter 2023) leverage principal component analysis to represent facial. However, this method falls short in capturing details such as wrinkles, the interior of the mouth, and hair, so it may not fully satisfy appearance requirements. i3DMM (Yenamandra et al. 2021) introduces an implicit function to model both the geometry and appearance of the human head with various attributes like shape, expression, and hairstyle. Both approaches rely on 3D data for their modeling. Recent advancements in Neural Radiance Fields (NeRF) (Mildenhall et al. 2020; Barron et al. 2021) have excelled in novel view synthesis due to their compact and powerful representation capability, relying solely on a set of multi-view images. Consequently, numerous works (Gafni et al. 2021; Park et al. 2021) have yielded detailed geometry (Zheng et al. 2022; Xu et al. 2023) and achieved a photo-realistic appearance (Zheng et al. 2023; Kirschstein et al. 2023) in modeling heads.
@@
Sign distance fields (SDF) (Park et al. 2019) and occupancy fields (Mescheder et al. 2019), showcase their representative ability over explicit representations, i.e. mesh, and point cloud. DVR (Niemeyer et al. 2020) and IDR (Yariv et al. 2020) focus on differentiating the surface rendering pipeline based on multi-view images. They incorporate corresponding masks to distinguish objects from the background during the training process. NeuS (Wang et al. 2021) and VolSDF (Yariv et al. 2021) refine the geometry reconstruction of NeRF by introducing a scheme that converts SDF to density, enhancing the surface representation in NeRF. Recent approaches (Rosu and Behnke 2023; Wang, Skorokhodov, and Wonka 2022, 2023) combine volume rendering with multi-scale hashing in Instant-NGP (Müller et al. 2022) and displacement fields to learn detailed surfaces through implicit functions from multi-view images.
@@
We summarize three categories primary approaches. The first involves an algorithm based on adversarial generation networks, aimed at deceiving unauthorized facial recognition by generating fake images highly similar to the original ones (Li, Wang, and Li 2019). However, this method is primarily targeted at public social platforms (Ciftci, Yuksek, and Demir 2023) or specific facial recognition (FR) systems that provide data for learning (Cherepanova et al. 2021). The second category involves the use of cryptographic techniques. Cryptographic techniques, including homomorphic encryption (Huang et al. 2020), secure multiparty computation (Ma et al. 2019), and other encryption primitives (Kou et al. 2021), are used to encrypt original images securely. These approaches introduce higher latency and computational costs. The third group of methods focuses on obfuscation techniques. These methods encompass actions such as implementing blurring (Jiang et al. 2023), introducing noise (Zhang et al. 2021), applying masking (Seneviratne et al. 2022), utilizing filtering (Zhou and Pun 2020), and employing image transformation (Wang, Kelly, and Veldhuis 2021). While practical, these methods irreversibly degrade image quality and may fail in subsequent tasks without a robust end-to-end solution. mm3DFace (Xie et al. 2023) uses mmWave signals for privacy protection, but its goal is not to reconstruct detailed geometry.
Method
Figure 1: Pipeline. We simulate the data flow process where our model reconstructs the human head geometry, without depending on sensitive images. Users can select photos that require privacy protection through a general operator denoted by $\sigma$. Following this, privacy-neutral images and privacy-protected images are used for geometric reconstruction. The reconstruction process occurs in two stages: stage 1 uses privacy-neutral images to establish the foundational geometry, while stage 2 employs image gradient information to refine the facial geometry further.
Preliminaries
Neural volume rendering. As demonstrated by NeRF (Mildenhall et al. 2020), it characterizes a 3D scene by employing volume density and color fields. Given known camera poses and ray directions, we conduct sampling along the rays, predicting both the color $\mathbf{c}_{i}$ and density $\sigma_{i}$ at each sample point $\mathbf{x}_{i}$ using Multi-Layer Perceptron (MLPs). The volume rendering process entails the integrating color radiance across the sampled points along the ray. Consequently, we approximate the rendered color for a specific pixel as:
@@
Recent contributions have been centered around neural volume rendering based on SDF representation (Wang et al. 2021; Yariv et al. 2021). These methods utilize MLPs to implicitly represent a 3D scene by predicting the SDF and color $\mathbf{c}_{i}$ relative to the viewpoint. This differs from the initial approach presented in (Mildenhall et al. 2020), which focused on predicting color and density. The extraction of the zero-level surface of the SDF in these methods results in a more reasonable geometric representation, showcasing significant effectiveness in reconstructing objects with smooth geometry.
@@
Two-stage Training
In our assumptions, the reconstruction process should neither access nor alter sensitive facial images, ensuring that the algorithm maintains practical utility for geometric reconstruction while fully adhering to privacy requirements. To implement this approach, all images used for reconstruction must be non-sensitive and are categorized into two types: privacy-neutral and privacy-protected. Neutral pictures, which do not contain facial information, such as those captured from the back of the head, are directly utilized in the training process. Privacy-protected images, initially taken from frontal perspectives, are processed through a specialized operator to ensure they meet our privacy goals before being uploaded by the user and used in the second stage.
In our proposed method, we employ a two-stage training framework for geometric reconstruction in Figure 1. In the first stage, we train a neural radiance field using neutral images, which enables the reconstruction of essential low-frequency information. The success of this stage depends on the quantity of privacy-neutral data users provide. While this information might have been deemed less valuable for facial geometry learning in the past, it proves to be beneficial in our method. Unless unavoidable circumstances require the use of templates, as discussed in the section Optimization, the privacy-neutral images provided by the user can be fully utilized. This enhances the geometric accuracy in the first stage of reconstruction, bringing it closer to the user’s geometric features, and ultimately contributing to the overall quality of the head reconstruction in the second stage.
In the second stage, we utilize privacy-protected data uploaded by users for the reconstruction process. Our supervision focuses solely on color variation information inherent in the original images. We intentionally refrain from utilizing full RGB information to mitigate the risk of privacy exposure. While the privacy-protected images appear visually unfriendly and blurry, they contain valuable color variation details crucial for refining geometric intricacies. Building on the foundation established in the first stage, we train for new frontal face viewpoints. We compute gradients of the rendered images, transforming them into multi-view color gradient modulus through a general operator $\sigma$. These modulus are then compared to the facial privacy-protected gradient information obtained from the user after passing through the operator. To optimize this stage, we introduce a novel loss equation, the gradient loss $\mathcal{L}_{\text{grad}}$, to guide the learning of geometric information in the second stage: